OpenStudy (anonymous):

Sam and Harry are family. Sam is currently is three times Harry's age. Sam's age is also 10 more than twice Henry's age. What are their current ages? A.) Sam is 24 years old, and Harry is 8 years old. B.) Sam is 39 years old, and Harry is 13 years old. C.) Sam is 27 years old, and Harry is 9 years old. D.) Sam is 30 years old, and Harry is 10 years old.

OpenStudy (xguardians):

D. Just apply this: Sam's age is also 10 more than twice Henry's age. until you find the correct answer, that is what I had done.

OpenStudy (anonymous):

Oh i get it thank you

OpenStudy (anonymous):

can you help me with another one please

OpenStudy (xguardians):

Sure create a new topic :)

OpenStudy (anonymous):

Maya had $27. She spent all the money on buying 3 burgers for $x each and 2 sandwiches for $y each. If Maya had bought 2 burgers and 1 sandwich, she would have been left with $11. A student concluded that the price of each burger is $5 and the price of each sandwich is $6. Which statement best justifies whether the student's conclusion is correct or incorrect? A.) The student's conclusion is correct because the solution to the system of equations 3x + 2y = 11 and 2x + y = 16 is (5, 6). B.)The student's conclusion is incorrect because the solution to the system of equations 3x – 2y = 11 and 2x – y = 16 is (5, 6). C.) The student's conclusion is correct because the solution to the system of equations 3x + 2y = 27 and 2x + y = 16 is (5, 6). D.) The student's conclusion is incorrect because the solution to the system of equations 2x + 3y = 27 and x + 2y = 16 is (5, 6).

OpenStudy (secret-ninja):

C